# Electrical spin manipulation in double SrTiO3/LaAlO3 quantum dots

### v1

@@ Abstract
The spin dynamics in two electron double quantum dots embedded in two dimensional electron gas at the interface between SrTiO3 and LaAlO3 is studied by an exact numerical solution of the time-dependent Schrödinger equation, in the context of the electric dipole spin resonance experiment. Based on the three band model of -electrons localized at Ti ions on the square lattice we analyze in details the singlet-triplet transition induced by the AC electric field, in the magnetic field range close to the avoided crossing which appears as a result of the spin-orbit coupling. Our calculations show that for symmetric double quantum dots the single photon spin-flip transitions is prohibited due to the parity symmetry and the transition can occur only by the higher order two-photon processes. For a weakly asymmetric system, when the first order singlet-triplet transitions are released due to the parity symmetry breaking, the spin-flip transition has a character of the Rabi oscillations for a low electric field amplitude. As the amplitude is increased the frequency of the transition is blueshifted (redshifted) for the magnetic field below (above) the single-triplet avoided crossing. Interestingly, for a sufficiently high magnetic field and high AC field amplitude the electric field drives the system across the avoided crossing inducing the spin-flip by the Landau-Zener-Stueckelberg-Majorana transitions with 100% spin flip probability for a slow sweep. Finally, the optimization of the geometrical parameters of the system with respect to the time of spin-flip of its fidelity is also presented.
@@
The ability to control the spin of electrons confined in quantum dots (QDs) has attracted growing interest in recent years, as it lies at the heart of developments towards a scalable spin-based quantum computer Burkard et al. (2023); Nadj-Perge et al. (2010a). The effiency of spin manipulation is entirely contingent upon the properties of the material platform, which dictate the strength of the spin-orbit coupling (SOC) Giglberger et al. (2007); Rashba and Efros (2003); Bychkov and Rashba (1984) or the hyperfine interaction with the nuclear spin bath. The latter is a source of spin decoherence Khaetskii et al. (2002) limiting the usability of a specific material platform in accordance with one of the Loss di Vincenzo criteria Loss and DiVincenzo (1998), whereby the quantum system must remain coherent for times much longer than the duration of elementary logic gates. For this reason, silicon-based QDs have been intensively studied over the last years due to their prolonged spin coherence time Mills et al. (2022).
@@
With respect to that, another promising but still unexplored platform for QD-based spin qubits technology are transition metal oxide heterostructures, such as SrTiO3/LaAlO3 (LAO/STO) interfaces Joshua et al. (2012); Maniv et al. (2015); Monteiro et al. (2019); Pavlenko et al. (2012); Biscaras et al. (2012). The electronic properties of the two-dimensional electron gas (2DEG) formed at the LAO/STO interface is determined by the 3-orbitals Diez et al. (2015); Khalsa et al. (2013); van Heeringen et al. (2013); Wójcik et al. (2021); Zegrodnik and Wójcik (2020), implying that direct and indirect decoherence stemming from interaction with the nuclear bath can be significantly mitigated, as it is proportional to the square of the wave function at the nuclei position Khaetskii et al. (2002) - this value diminishes to zero for d-electrons. This property, combined with the strong spin-orbit interaction Caviglia et al. (2010); Yin et al. (2020), high mobility Ohtomo and Hwang (2004), and susceptibility to electrostatic gates at a level comparable to semiconducting materials Bjørlig et al. (2020); Thierschmann et al. (2018); Guenevere et al. (2017), initiates the currently ongoing research on LAO/STO-based quantum dots as potential spin-qubits with inherent scalability of 2D systems. The first experimental realization of electrostatically defined LAO/STO QDs has been already reported, yielding a Coulomb blockade diamond characteristic for a well-defined confinement Bjørlig et al. (2020); Thierschmann et al. (2018).
@@
The demonstration of electrical spin manipulation in these systems constitutes the next milestone in spin-qubit realization based on the transition metal oxide heterostructures. Significant progress in this field has been recently achieved in electrostatically defined semiconducting QDs Schroer et al. (2011), where the microwave magnetic field, previously used to control spin dynamics Koppens et al. (2006, 2008), has been replaced by the ac gate voltage Nadj-Perge et al. (2010a, b, 2012); Golovach et al. (2006). In this technique, called electric dipole spin resonance (EDSR), the microwave potential applied to one of the gates generates a time-dependent potential which, due to the momentum-dependent spin-orbit interaction Giglberger et al. (2007); Rashba and Efros (2003); Bychkov and Rashba (1984), hybridizes the spin-up and spin-down states, leading to Rabi oscillations between the Zeeman-split electronic states Koppens et al. (2006). For two electrons, the SOC hybridization induces level repulsion between the spin singlet and triplet states, which is often used to determine the strength and direction of the effective SOC field Stepanenko et al. (2012). Recently, EDSR technique applied to double quantum dots (DQDs) allowed for a demonstration of Pauli spin blockade Nowack et al. (2007); Nadj-Perge et al. (2010a). In this regime, the system initially set to the spin triplet state , with tunneling from the left to the right QDs prohibited by selection rules, transmits to the spin singlet state , which is possible due to the SOC generated by the ac field applied to one of the gates Nowak et al. (2012).
@@
The manuscript is organized as follows: in Sec. II we present a theoretical model of double quantum dots in 2DEG at the (001) LAO/STO interface as well as the scheme used for time-dependent two electron calculations. Sec. III contains the analysis of electronic spectrum of a two-electron double quantum dots as well as results of time dependent simulations (EDSR), finally summary and conclusions are included in Sec. IV.
@@ II.1 The single-electron problem
We consider 2DEG in the quantum well created at the (001) LAO/STO interface Popović et al. (2008); Pavlenko et al. (2012). The system is characterized by a significant atomic SO coupling and a strong Rashba interaction due to the asymmetry of the vertical electric field at the interface Diez et al. (2015); van Heeringen et al. (2013); Khalsa et al. (2013). We use the real space representation Szafran et al. (2023) of the tight-binding Hamiltonian Diez et al. (2015) spanned by orbitals of Ti ions that are arranged in a square lattice. The Hamiltonian of the sytem is given by
@@
where the matrix corresponds to with the orbital angular momentum represented in the -orbitals basis. Finally, the coupling of the external magnetic field to the spin and atomic orbital momentum of electrons is taken into account by the Hamiltonian
where is the Bohr magneton, is the Landé factor, with and with
@@
Finally, the hybridization between , orbitals are taken into account by
In our simulations, we take on the tight-binding parameters meV, meV, meV, meV, after Ref. Zegrodnik and Wójcik (2020); Diez et al. (2015), the Landé factor Ruhman et al. (2014); Bjørlig et al. (2020) and the SO coupling parameters meV, meV Caviglia et al. (2010); Yin et al. (2020).
The external magnetic field induces the Aharonov-Bohm phase shifts for the electrons moving along the interface, that we account for using the Peierls approach, i.e. multiplying the elements of the Hamiltonian corresponding to the hopping between ions placed at points and by a factor , where is the vector potential taken in the symmetric form for .
@@
where is the distance between the minima of the separate functions (centers of the left and right QDs), and define the depth of each potential minimum and is the size parameter of the separate Gaussians.
@@ II.2 The two-electron problem
with the dielectric constant Gariglio et al. (2015). The problem is solved with the exact diagonalization method with the basis spanned by the atomic orbitals
@@ III.1 The two-electron eigenstates
For a symmetric potential the single-electron Hamiltonian eigenstates are also eigenstates parity operator , where is the scalar parity operator . The lowest-energy singlet and the lowest spin-polarized triplet that participate in the avoided crossing have the same – negative – parity – see Fig. 4. The same symmetry of the lowest singlet and triplet states has a direct consequence for the spin transitions driven by AC electric field, namely the dipole transition matrix elements between the two lowest-energy levels vanish so the first-order (single-photon) transitions is forbidden and only the second-order processes (two-photon) can occur. Similar behavior has been observed for two electrons in single quantum dot and discussed in details in our previous paper Szafran et al. (2023).
@@ III.2 EDSR in weakly asymmetric system
Fig. 5(a) shows the spectrum for nm near the singlet-triplet avoided crossing that is shifted to higher magnetic field by the asymmetry due to the increased exchange energy [Fig. 3(a)] (to T from T in the case). The width of the avoided crossing is also enlarged: from 12.3eV (for meV) to 13.8 eV (for meV). The essential quantity for the first-order spin transitions is the dipole matrix element between the two-lowest energy levels which is plotted by the black curve in Fig. 5(b). For an ideally symmetric system a similar singlet-triplet avoided crossing of energy levels is obtained with the interchange of the average values of the spin between the two lowest-energy levels, but the transition dipole matrix element is zero. For the system with lifted parity symmetry, the matrix element reaches its maximum in the center of the avoided crossing which allows the first order transition possible.
@@
For the magnetic fields in Fig. 6(a,b) and Fig. 6(c,d) the ground-state is singlet and triplet, respectively. At the amplitude of 5 eV the largest peak corresponds to the singlet-triplet transition that occurs at the driving energy near the singlet-triplet energy difference. For T and T, where the dipole matrix element is large, at the lowest-amplitude, we also resolve the peaks of higher-order transitions at half and rd of the frequency for the single-photon transition processes. As the AC field amplitude is increased the frequency of the spin-flip transition [Fig. 6(a,b)] for the magnetic field below the single-triplet avoided crossing is blueshifted at higher AC field amplitudes. On the other hand the frequency for the spin transition for the magnetic field above the single-triplet avoided crossing gets redshifted for high AC amplitudes [Fig. 6(c,d)]. In both cases, in an intense AC field, we find an effective upbeat of the triplet energy with respect to the singlet. In Fig. 6(c) we find that the regular transition spectrum gets deteriorated at higher frequency with the disappearance of the single-photon peak. For T at higher amplitude the AC field drives the system across the singlet-triplet avoided crossing so that the process for the spin-flip acquires the Landau-Zener-Stueckelberg-Majorana (LZSM) character Ivakhnenko et al. (2023).
@@
Fig. 7 shows the two lowest-energy states for the double dot system with a time-independent electric field of added to the potential. The avoided crossing for T [Fig. 7(b)] is reached at V. Note, that in AC field with the amplitude of V/nm the maximal occupancy of the singlet-state is no longer zero at low limit [cf. Fig. 6(c)]. At higher amplitude the low-frequency limit in Fig. 6(c) corresponds to an adiabatic sweep of the potential across the singlet-triplet avoided crossing with 100% spin flip probability for a slow sweep. For nonzero AC frequency the LZSM interference between the contributions from both the states appear with a diabatic and an adiabatic components Ivakhnenko et al. (2023) generating the interference comb visible in Fig. 6(c) at higher amplitude and low frequency. For T we reach a similar point at V/nm [cf. Fig. 6(d) and Fig. 7(c)]. The change of the character of the triplet-singlet transition is better illustrated by the transfer probability plotted as a inverse of the driving energy presented in Fig. 8 for T. As the amplitude increases from 5V/nm to 25V/nm higher-order peaks corresponding to multiphoton transitions appear but the corresponding peaks fall below 100% at large . For V/nm and above, the transition probability peaks acquire the height of 100% that does not fall below at large , and appear periodically on the scale, which are characteristics of the LZSM interference Ivakhnenko et al. (2023).
@@
Figure 9(a) shows the position of the first order transition on the energy scale as a function of the AC amplitude. Note that, the magnetic fields considered in Figs. 9(a) as well as 6 were chosen to produce equal singlet-triplet energy splitting before and after the avoided crossings, hence the coalescent lines at the limit of zero amplitude. The frequency shifts before and after the singlet-triplet avoided crossing have opposite signs but similar magnitudes. The line for T is interrupted above 35 V/nm, since the resonance line can be hardly recognized in the spectrum [cf. Fig. 6(c)].

### v2

@@ Abstract
The spin dynamics in two electron double quantum dots embedded in two dimensional electron gas at the interface between SrTiO3 and LaAlO3 is studied by an exact numerical solution of the time-dependent Schrödinger equation in the context of the electric dipole spin resonance experiment. Based on the three band model of -electrons localized at Ti ions on the square lattice we analyze in details the singlet-triplet transition induced by the AC electric field, in the magnetic field range close to the avoided crossing which appears as a result of the spin-orbit coupling. For symmetric double quantum dots the single photon spin-flip transition is prohibited due to the parity symmetry and the transition can occur only by the higher order two-photon processes. For a weakly asymmetric system, when the first order singlet-triplet transitions are released due to the parity symmetry breaking, the spin-flip transition has a character of the Rabi oscillations for a low electric field amplitude. As the amplitude is increased the frequency of the transition is blueshifted (redshifted) for the magnetic field below (above) the single-triplet avoided crossing. For a sufficiently high magnetic field and high AC field amplitude the electric field drives the system across the avoided crossing inducing the spin-flip by the Landau-Zener-Stueckelberg-Majorana effect. Finally, the optimization of the geometrical parameters of the system with respect to the time of spin-flip of its fidelity is also presented.
@@
The control the spin of electrons confined in quantum dots (QDs) has attracted a growing interest in recent years, as it lies at the heart of developments towards a scalable spin-based quantum computer Burkard et al. (2023); Nadj-Perge et al. (2010). The efficiency of spin manipulation is entirely contingent upon the properties of the material platform, which dictate the strength of the spin-orbit coupling (SOC) Giglberger et al. (2007); Rashba and Efros (2003); Bychkov and Rashba (1984) or the hyperfine interaction with the nuclear spin bath. The latter is a source of spin decoherence Khaetskii et al. (2002) limiting the usability of a specific material platform in accordance with one of the Loss di Vincenzo criteria Loss and DiVincenzo (1998), whereby the quantum system must remain coherent for times much longer than the duration of elementary logic gates. For this reason, silicon-based QDs have been intensively studied over the last years due to their prolonged spin coherence time Mills et al. (2022).
@@
With respect to that, another promising but still unexplored platform for QD-based spin qubits technology are transition metal oxide heterostructures, such as SrTiO3/LaAlO3 (LAO/STO) interfaces Joshua et al. (2012); Maniv et al. (2015); Monteiro et al. (2019); Pavlenko et al. (2012); Biscaras et al. (2012). The electronic properties of the two-dimensional electron gas (2DEG) formed at the LAO/STO interface is determined by the 3-orbitals Diez et al. (2015); Khalsa et al. (2013); van Heeringen et al. (2013); Wójcik et al. (2021); Zegrodnik and Wójcik (2020), implying that direct and indirect decoherence stemming from interaction with the nuclear bath can be significantly mitigated, as it is proportional to the square of the wave function at the nuclei position Khaetskii et al. (2002) - this value diminishes to zero for d-electrons. This property, combined with the strong spin-orbit interaction Caviglia et al. (2010); Yin et al. (2020), high mobility Ohtomo and Hwang (2004), and susceptibility to electrostatic gates at a level comparable to semiconducting materials Bjørlig et al. (2020); Thierschmann et al. (2018); Guenevere et al. (2017), initiates the currently ongoing research on LAO/STO-based quantum dots as potential spin-qubits with inherent scalability of 2D systems. The first experimental realization of electrostatically defined LAO/STO QDs has been already reported, yielding a Coulomb blockade diamond characteristic for a well-defined confinement Bjørlig et al. (2020); Thierschmann et al. (2018). The lateral confinement within LAO/STO interface was created earlier with the litography techniques Cen et al. (2009) and a detection of the confined charge island, or a quantum dot in a system tailored with nanolitography has been reported Maniv et al. (2016).
@@
The demonstration of electrical spin manipulation in these systems constitutes the next milestone in spin-qubit realization based on the transition metal oxide heterostructures. Significant progress in this field has been recently achieved in electrostatically defined semiconducting QDs Schroer et al. (2011), where the microwave magnetic field, previously used to control spin dynamics Koppens et al. (2006, 2008), has been replaced by the AC gate voltage Nadj-Perge et al. (2010, 2010, 2012); Golovach et al. (2006). In this technique, called electric dipole spin resonance (EDSR), the microwave potential applied to one of the gates generates a time-dependent potential which, due to the momentum-dependent spin-orbit interaction Giglberger et al. (2007); Rashba and Efros (2003); Bychkov and Rashba (1984) and the resulting effective magnetic field drives transitions between hybrydized spin-up and spin-down states, leading to Rabi oscillations Koppens et al. (2006). For two electrons, the SOC hybridization induces level repulsion between the spin singlet and triplet states, which is often used to determine the strength and direction of the effective SOC field Stepanenko et al. (2012). Recently, EDSR technique applied to double quantum dots (DQDs) allowed for a demonstration of Pauli spin blockade Nowack et al. (2007); Nadj-Perge et al. (2010). In this regime, the system initially set to the spin triplet state , with tunneling from the left to the right QDs prohibited by selection rules, transmits to the spin singlet state , which is possible due to the SOC generated by the AC field applied to one of the gates Nowak et al. (2012).
@@
The present paper deals only with the electronic part of the system without the phonon-field. The coupling of electrons with the charge environment via acoustic phonons, that leads to the effects of relaxation and decoherence Fischer et al. (2009); Climente et al. (2007), needs to be integrated in a future model to account for these effects. The phonon field at the LAO/STO interface is a subject of only recent studies, in particular the effects of the surface on the optical phonon branch near the energy of meV has been determined in Ref. Liu et al. (2023).
The manuscript is organized as follows: in Sec. II we present a theoretical model of double quantum dots in 2DEG at the (001) LAO/STO interface as well as the scheme used for time-dependent two electron calculations. Sec. III contains the analysis of electronic spectrum of two-electron double quantum dots as well as results of time dependent simulations (EDSR), finally summary and conclusions are included in Sec. IV.
@@ II.1 The single-electron problem
We consider 2DEG in the quantum well created at the (001) LAO/STO interface Popović et al. (2008); Pavlenko et al. (2012). The system is characterized by a significant atomic SO coupling and a strong Rashba interaction due to the asymmetry of the vertical electric field at the interface Diez et al. (2015); van Heeringen et al. (2013); Khalsa et al. (2013). We use the real space representation Szafran et al. (2024) of the tight-binding Hamiltonian Diez et al. (2015) spanned by orbitals of Ti ions that are arranged in a square lattice. The Hamiltonian of the system is given by
@@
where the tight-binding parameters meV, meV, meV, are adopted after Ref. Zegrodnik and Wójcik (2020); Diez et al. (2015), and is the external potential given at the ion position (.)
@@
where the matrix corresponds to with the orbital angular momentum represented in the -orbitals basis and we take the SO coupling parameter meV Caviglia et al. (2010).
Finally, the coupling of the external magnetic field to the spin and atomic orbital momentum of electrons is taken into account by the Hamiltonian
where is the the Landé factor Ruhman et al. (2014); Bjørlig et al. (2020), is the Bohr magneton, with and with
@@
with the Rashba SO coupling parameter meV Yin et al. (2020). Finally, the hybridization between , orbitals are taken into account by
with the hybridization parameter meV Diez et al. (2015).
The external magnetic field induces the Aharonov-Bohm phase shifts for the electrons moving along the interface, that we account for using the Peierls approach, i.e. multiplying the elements of the Hamiltonian corresponding to the hopping between ions placed at points and by a factor , where is the vector potential taken in the symmetric form for . Most of the results provided below are obtained for this perpendicular magnetic field orientation. However, the results for in-plane orientation are also provided. The in-plane field does not induce any Aharonov-Bohm phase shifts, produces no orbital effects and enters to the Hamiltonian only via the spin Zeeman term.
@@
where is the distance between the minima of the separate functions (centers of the left and right QDs), and define the depth of each potential minimum and nm is the size parameter of the separate Gaussians.
In our model we assume that the potential of each of the dots separately is rotationally symmetric. For the parameters applied below the tunnel coupling between the dots is non-negligible, so that the single-electron wave functions are extended over both the dots. Any deformation of the Gaussian potential with one of the dots will predominantly act as an effective potential shift in one of the dots. The shift would induce an asymmetry in the double dot potential that, as we show below, has a very pronounced effect on the dipole matrix elements.
@@ II.2 The two-electron problem
with the dielectric constant Gariglio et al. (2015). The dielectric constant of STO can reach the values as large as 10 000 but can be limited in an electric field. As shown in the Appendix to Ref. Szafran et al. (2024) the electric field near the interface hosting 2DEG is very strong and leads to a reduction of the value to the one we apply in the present paper.
The problem is solved with the exact diagonalization method with the basis spanned by the atomic orbitals
@@ II.3 Time stepping
The present model does not take into account the relaxation and dephasing mechanisms. In a paper by two of the present authors Nowak and Szafran (2014) we considered the effects of the relaxation via acoustic phonons for EDSR experiments involving two states and demonstrated that the effect may lead to an off-resonant transition to a third state with a lower energy than the two states coupled through a resonant AC pulse. In Ref. Szafran et al. (2024) we argued that, when the considered transitions involve the two-lowest energy states in the spectrum, the relaxation mechanism is of a secondary importance. The dephasing in III-V and Si quantum dots is mainly due to the hyperfine interactions with the nuclear spins which should be absent in STO materials since the electron states are spanned by the 3d orbitals. The dephasing due to the interaction with the charge disorder in the surrounding matrix should be of a limited importance given the large value of the dielectric constant.
@@ III.1 The two-electron eigenstates
For a symmetric potential the single-electron Hamiltonian eigenstates are also eigenstates of the parity operator , where is the scalar parity operator . The lowest-energy singlet and the lowest spin-polarized triplet that participate in the avoided crossing have the same – negative – parity – see Fig. 4. The same symmetry of the lowest singlet and triplet states has a direct consequence for the spin transitions driven by AC electric field, namely the dipole transition matrix elements between the two lowest-energy levels vanish so the first-order (single-photon) transitions is forbidden and only the second-order processes (two-photon) can occur. Similar behavior has been observed for two electrons in single quantum dot and discussed in details in our previous paper Szafran et al. (2024).
@@ III.2 EDSR in weakly asymmetric system
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | 2.58 | 2.54 | 0.45 | 0.4 |
2 | 0.041 | 0.23 | 0.31 | 0.15 |
3 | ||||
4 | ||||
5 | ||||
6 |
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | 0 | 0 | 0 | 0 |
2 | 0.138 | 0.051 | 0.051 | 0.138 |
3 | 0.213 | 0.212 | 0.261 | 0.346 |
4 | 0.292 | 0.374 | 0.182 | 0.689 |
5 | 1.318 | 1.317 | 1.362 | 1.449 |
6 | 2.640 | 2.639 | 2.684 | 2.772 |
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | -1.94 | -1.98 | ||
2 | -1.99 | -1.96 | ||
3 | ||||
4 | 1.985 | 1.985 | 1.985 | 1.985 |
5 | ||||
6 |
Fig. 5(a) shows the spectrum for nm near the singlet-triplet avoided crossing that is shifted to higher magnetic field by the asymmetry due to the increased exchange energy [Fig. 3(a)] (to T from T in the case). The width of the avoided crossing is also enlarged: from 12.3eV (for meV) to 13.8 eV (for meV). The essential quantity for the first-order spin transitions is the dipole matrix element between the two-lowest energy levels which is plotted by the black curve in Fig. 5(b) and listed in Table I for selected values of the magnetic field. In Table I the first row corresponds to the average value of for the ground-state which are about 2.5 nm due to the shift of the wave function to the deeper dot. The shift gets smaller after the avoided crossing when the spin-polarized triplet replaces the spin-singlet in the ground-state. By Pauli exclusion principle the electrons in the triplet state cannot contain an admixture of configuration with both electrons in the single-electron ground-state of the deeper dot, hence the smaller shift to the deeper quantum dot for the triplet state. For an ideally symmetric system a similar singlet-triplet avoided crossing of energy levels is obtained with the interchange of the average values of the spin between the two lowest-energy levels, but the transition dipole matrix element is zero. For the system with lifted parity symmetry, the matrix element reaches its maximum in the center of the avoided crossing which allows the first order transition possible.
@@
Fig. 6 shows the maximal probability of the singlet-triplet transitions that are found during 5 ns simulation for a range of AC amplitudes and four selected perpendicular magnetic fields 0.434 T, 0.94 T, 1.5 T and 2 T. The matrix dipole elements for the singlet-triplet transitions, the energy above the ground-state and the z-component of the total spin for the four selected magnetic field values are given in Table I, II and III, respectively. The values are selected symmetrically with respect to the singlet-triplet avoided crossing – see second row of Table II. For the magnetic fields in Fig. 6(a,b) and Fig. 6(c,d) the ground-state is singlet and triplet, respectively. At the amplitude of 5 eV the largest peak corresponds to the singlet-triplet transition that occurs at the driving energy near the singlet-triplet energy difference. For T and T, where the dipole matrix element is large, at the lowest-amplitude, we also resolve the peaks of higher-order transitions at half and rd of the frequency for the single-photon transition processes. In an ideal symmetric system that we considered previously Szafran et al. (2024) the first-order singlet-triplet transitions are forbidden by the parity symmetry reasons. The transition matrix element vanish for the states that have the same parity as in the case of the singlet and the spin-polarized triplet in an inversion-symmetric system (cf. Fig. 4). The singlet-triplet transitions can still be induced for a symmetric system but with a subharmonic, two-photon transitions. The two-photon transitions at low AC frequency appear at the energy equal to roughly half the energy difference between the singlet and triplet states. The multiphoton transitions are higher order processes that require a much longer spin-flip time than the single-photon transitions. As the AC field amplitude is increased the frequency of the spin-flip transition [Fig. 6(a,b)] for the magnetic field below the single-triplet avoided crossing is blueshifted at higher AC field amplitudes. On the other hand the frequency for the spin transition for the magnetic field above the single-triplet avoided crossing gets redshifted for high AC amplitudes [Fig. 6(c,d)]. In both cases, in an intense AC field, we find an effective upbeat of the triplet energy with respect to the singlet.
Fig. 7 shows the time-resolved occupation of the ground-state singlet (black line) and the excited triplet (red line) as well as the overal occupation of higher excited states (blue) line for T for the amplitude of the AC field increasing from 10eV/nm (the highest row of plots) to 40eV/nm (the lowest row of plots). The evolution shows the results corresponding to Fig. 6(a), only longer evolution times than 5ns were taken when necessary. The left column of the plots in Fig. 7 shows the time evolution for the resonant single-photon frequency and the right column of the plots to the first subharmonic (two-photon) transition. The AC frequencies corresponds to the rightmost and second to the right peaks in Fig. 6(a). Note the difference of the times necessary to flip the spins in the single- and two-photon-transitions (left vs right column of the plots in Fig. 7). The blue line in Fig. 7 corresponds to the overall contributions of the higher energy-states. This contribution is nearly missing for the lowest amplitude considered in Fig. 7(a,b), and grows for higher AC amplitudes. Note, that the triplet occupation (the red line in Fig. 7) preserves its sine form also for higher AC amplitudes. On the other hand there is a fast oscillation that is superposed on the cosine form of the occupation line for the ground-state singlet. The ground-state singlet is strongly dipole-coupled to the 5th and 6th singlet states (see Table I) that are much higher in energy (Table II) with the large energy difference leading to a fast rate of these off-resonant transitions between the ground-state and excited singlets. The contribution of the higher excited states vanishes when the occupation of the ground-state singlet is zero. Figure 8 shows the overall contribution for all the excited states (black line) in the time evolution of Fig. 7(g) and the separate contributions from the 5-th and 6-th states (black line and blue line, respectively). The contribution of the 5-th state is larger consistently with the values of the dipole matrix element of Table I (see the first column of Table I corresponding to the perpendicular magnetic field of 0.434 T). Contribution of higher energy states is negligible.
The spectrum for the magnetic field increased from 0.434T to 0.94T [Fig. 6(b)] has much better resolved subharmonic structure due to times larger dipole matrix element (the second row of Table I and Fig. 5(b)). In Fig. 6(c) for 1.5T we find that the regular transition spectrum gets deteriorated at higher frequency with the disappearance of the single-photon peak. A near structure of peaks is restored at 2T – see Fig. 6(d). Both Fig. 6(c,d) in contrast to Fig. 6(a,b) develop a large values of spin-flip probability at the lowest range of the frequency. This is due to the electric field sweeping across the singlet-triplet avoided crossing. In order to explain this effect in Fig. 9 we show the two lowest-energy states for the double dot system with a time-independent electric field of added to the potential. The avoided crossing for T [Fig. 9(b)] is reached at V. Note, that in AC field with the amplitude of V/nm the maximal occupancy of the singlet-state is no longer zero at low limit [cf. Fig. 6(c)]. At higher amplitude the low-frequency limit in Fig. 6(c) corresponds to an adiabatic sweep of the potential across the singlet-triplet avoided crossing with 100% spin flip probability in a limit of an adiabatic sweep.
@@
For T and T at higher amplitude the AC field drives the system across the singlet-triplet avoided crossing so that the process for the spin-flip can acquire the Landau-Zener-Stueckelberg-Majorana (LZSM) character Ivakhnenko et al. (2023). The ground-state system that is driven across the avoided crossing can either stay in the ground-state (adiabatic limit) with the transformation of the wave function from the one characteristic to the ground-state at the other side of the avoided crossing or preserve its wave function and find itself in the excited state. For periodic modulation the phases for both processes accumulate or cancel that leads to the LZSM interference pattern Ivakhnenko et al. (2023). The interference produces the comb visible in Fig. 6(c) at higher amplitude and low frequency. For T we reach a similar point at V/nm [cf. Fig. 6(d) and Fig. 9(c)]. The change of the character of the triplet-singlet transition is better illustrated by the transfer probability plotted as a inverse of the driving energy presented in Fig. 10 for T. As the amplitude increases from 5V/nm to 25V/nm higher-order peaks corresponding to multiphoton transitions appear but the corresponding peaks fall below 100% at large . For V/nm and above, the transition probability peaks acquire the height of 100% that does not fall below at large , and appear periodically on the scale, which are characteristics of the LZSM interference Ivakhnenko et al. (2023).
The time-resolved occupation of the states involved in the system dynamics for the magnetic field of 2T is illustrated in Figs. 11 and 12. Fig. 11 corresponds to the one- and two-photon transitions. At the AC amplitude of V/nm [Fig. 11(a,b)] the occupation has a similar character to the ones found for T [cf. 7(a,b)]. At higher amplitude the contribution from the higher excited singlet states (blue line in Fig. 11(e-h)) is correlated with the red line – that now corresponds to the spin-singlet which is the first excited state in the field of 2T. For the amplitude of 20V/nm the system is still not driven across the avoided crossing, only approach it (see Fig. 9(c)). When this happens (cf. Fig. 11(e,f) and (g,h) for the amplitude of the AC field of 30 and 40V/nm, respectively) the time evolution of the occupations deviates visibly from the simplest periodicity – see the shapes of the maximal occupation probability at each turn. Note, that the time evolution in Fig. 7 for the magnetic field of 0.434 T, when the sweep across the avoided crossing does not occur, is periodic in the entire range of frequencies considered. In Fig. 11 for higher AC amplitudes the second order transition becomes nearly as fast as for the two-photon transition which never occurs for the driving off the avoided crossing. The driving in Fig. 11 corresponds to the first two peaks at right of Fig. 6(d), i.e. to a relatively large AC frequency. The time evolution for the lowest-frequency i.e., rightmost peak in Fig. 10 is given in Fig. 12(a). The orange line in this peak shows the curve that modulates the external field. The center of the AC of Fig. 9(c) is reached at . Above this value the occupation of the ground (black line) and first excited (red line) state in Fig. 12(a) is nearly constant. Variation of the occupation is found when the function acquires values lower than . In Fig. 12(a) transitions from the ground to the excited state after each sweep of the AC field accumulate which results in a constructive LZSM interference. For contrast, in Fig. 12(b) we took the driving frequency between the two rightmost peaks of Fig. 10. Here the effects of the transitions do not accumulate but cancel at each sweep.
@@
Figure 13(a) shows the position of the first order transition on the energy scale as a function of the AC amplitude. Note that, the magnetic fields considered in Figs. 13(a) and Fig. 6 were chosen to produce equal singlet-triplet energy splitting before and after the avoided crossings, hence the coalescent lines at the limit of zero amplitude. The frequency shifts before and after the singlet-triplet avoided crossing have opposite signs but similar magnitudes. The line for T is interrupted above 35 V/nm, since the resonance line can be hardly recognized in the spectrum [cf. Fig. 6(c)].